The sum of Jim's age and Mike's age is 52. Six years ago mike was four times as old as jim was. How old is each now?
Let x = jims age
let y = mikes age

Answer by mananth(16949)

The sum of Jim's age and Mike's age is 52. Six years ago mike was four times as old as jim was. How old is each now?
Let x = jims age
let y = mikes age
1...x+y=52
15 years ago their ages were x-6 & y-6
y-6=4(x-6)
y-6=4x-24
2...y-4x=-18
Subtract eq2 fro eq1
..
x+y-y+4x=52-(-18)
5x=60
x= 12 years
y= 30 years